Bus A runs every 8 minutes, Bus B runs every 10 minutes, and Bus C runs every 12 minutes. If all three buses leave the station at 8:00 a.m., when is the next time all three buses leave the station at the same time?
a.9:00 am
b.10:00 am
c.11:00 am
d.12:00 pm
e.1:00 pm

Found 2 solutions by stanbon, galactus:
Answer by stanbon(75887)

Bus A runs every 8 minutes, Bus B runs every 10 minutes, and Bus C runs every 12 minutes. If all three buses leave the station at 8:00 a.m., when is the next time all three buses leave the station at the same time?
---
Find the least common multiple of 8,10,12.
---
8 = 2^3
10 = 2*5
12 = 3*2^2
---
lcm = 2^3*3*5 = 120
---
120 min = 2 hrs.
-----
Ans: 10:00 am
===============
Cheers,
Stan H.
================
a.9:00 am
b.10:00 am
c.11:00 am
d.12:00 pm
e.1:00 pm

Answer by galactus(183)

What is the least common multiple of 8,10, and 12?.
It is 120.
So, the buses will be in sync in another 120 minutes or 2 hours later.
At 10 am